OpenStudy (anonymous):

Quick Algebra 2 question. Anybody??

OpenStudy (anonymous):

@kelliegirl33 @Loser66 @wyattjohnson @radar ?

OpenStudy (loser66):

call me and no question pop up?

OpenStudy (anonymous):

I'm doing an exponential growth assignment: A fisherman decided to stock his favorite lake with a few fish that are not native to the region. The fish he introduces have no natural predators in the lake and they experience a large exponential growth. The introduction of an invasive species can be very detrimental to an ecosystem and there have already been noted effects. The National Fish and Wildlife Foundation has asked for your help. I need help with calculating the population of the fish reaches 500.

OpenStudy (anonymous):

For p(x) I have 100,200,400,800

OpenStudy (loser66):

do you have exponential growth formula ?? P(x) = A *....something^ t

OpenStudy (phi):

Loser helped you on this 3 days ago.

OpenStudy (anonymous):

p(x)=100*2^x

OpenStudy (anonymous):

not this particular question @phi

OpenStudy (phi):

you want to find x when p(x) is 500 100 * 2^x = 500 what is the first step to solving this ?

OpenStudy (anonymous):

I'm guessing find x?

OpenStudy (phi):

to find x, first divide both sides by 100

OpenStudy (phi):

you can estimate when you get to 500 by looking at your table. to get the exact number, you solve 100 * 2^x =500 2^x = 5 take the log of both sides x * log(2) = log(5) divide both sides by log(2) x= log(5)/log(2) you need a calculator to find x.

OpenStudy (anonymous):

I can figure that one out from here. But I'm really having trouble with: The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm. I don't understand this at all

OpenStudy (phi):

they want you to start with your equation p(x)=100*2^x I think they want you to take the log base 2 (though it is easier to cut to the chase and take the log base 10). take the log base 2 of both sides \[ \log_2(p(x)) = \log_2(100 \cdot 2^x) \]

OpenStudy (phi):

then they want you to change the log base 2 to log base 10, using the "change of base" formula. \[ \log_2(x) = \frac{\log_{10}(x) }{\log_{10}(2)} \]